### v1

# Unpacking Positional Encoding in Transformers: A Spectral Analysis of Content-Position Coupling

Abstract
Positional encoding (PE) is essential for enabling Transformers to model sequential structure. However, the mechanisms by which different PE schemes couple token content and positional information—and how these mechanisms influence model dynamics—remain theoretically underexplored. In this work, we present a unified framework that analyzes PE through the spectral properties of Toeplitz and related matrices derived from attention logits. We show that multiplicative content-position coupling—exemplified by Rotary Positional Encoding (RoPE) via a Hadamard product with a Toeplitz matrix—induces spectral contraction, which theoretically improves optimization stability and efficiency. Guided by this theory, we construct synthetic tasks that contrast content-position dependent and content-position independent settings, and evaluate a range of PE methods. Our experiments reveal strong alignment with theory: RoPE consistently outperforms other methods on position-sensitive tasks and induces “single-head deposit” patterns in early layers, indicating localized positional processing. Further analyses show that modifying the method and timing of PE coupling, such as MLA in Deepseek-V3, can effectively mitigate this concentration. These results establish explicit content-relative mixing with relative-position Toeplitz signals as a key principle for effective PE design and provide new insight into how positional structure is integrated in Transformer architectures.
1 Introduction
Transformer architectures have become the prevailing paradigm for sequence modeling across diverse domains owing to their powerful capacity to capture long-range dependencies. A cornerstone of their success is the attention mechanism’s ability to capture dependencies regardless of token distance. However, lacking inherent order sensitivity, Transformers critically rely on positional encoding (PE) to inject information about token positions and enable the modeling of sequential structure vaswani2017attention . The field has seen a proliferation of diverse PE variants, ranging from simple additive and learnable absolute embeddings to relative biases raffel2020exploring and sophisticated multiplicative schemes like Rotary Positional Encoding (RoPE) SU2024127063 . Yet, despite extensive empirical evaluation comparing their performance on downstream tasks, our fundamental understanding of the internal mechanisms by which these different PE variants specifically couple content and positional information, and how this coupling influences the model’s dynamics and internal learned structure, remains limited.
Existing evaluations of PE methods often rely on aggregate metrics such as perplexity HuaFoPE2025 or on length-extrapolation benchmarks kazemnejad2023impact . While providing valuable performance comparisons, these benchmarks shed little light on how positional signals are actually processed or on the mechanisms by which different encodings influence model behavior. Moreover, a focus on length-generalization often implies that, within a fixed sequence length, all PE schemes perform similarly. However, to our knowledge, no prior work has systematically confirmed or challenged this perception regarding within-length performance differences. This lack of understanding and empirical investigation highlights a critical gap in developing a unified theoretical framework that rigorously connects the specific design principles of PE (i.e., how they formulate positional coupling) to their impact on the attention dynamics and learned internal structure.
In this paper, we bridge this gap by introducing a spectral-theoretic framework complemented by targeted empirical studies. As illustrated in Figure 1, our method begins with the decomposition of each token embedding into content and position components ($c_{i}$ and $p_{i}$). We then assume that the position component gives rise to a Toeplitz-structured term in the attention logits, and we characterize how various PE schemes—whether additive biases or multiplicative rotations—map to distinct Toeplitz (or related) matrix operations within the attention computation.
Figure 1: Schematic Overview. (Left) Our framework analyzes token decomposition and how PE mechanisms (additive, e.g., adding Toeplitz $\mathbf{B}$; multiplicative, e.g., Hadamard with relative-position Toeplitz $G_{\mathbf{e}}$) structure attention logits. (Right) Multiplicative coupling theoretically induces spectral contraction (tighter eigenvalue bounds), predicting better learning/generalization and single-head deposit patterns. Additive mechanisms show smaller gains and more diffuse positional processing.
Our framework shows that content–position coupling can take two principal forms in the attention logits: (i) additive mechanisms, where position information (e.g., relative-position encodings) contributes by adding an explicit Toeplitz matrix $\mathbf{B}$ to the logits; and (ii) multiplicative schemes, most notably RoPE’s mechanism involving a Hadamard product with the relative-position Toeplitz matrix $G_{\mathbf{e}}$. Applying spectral analysis techniques, including Szegő’s theorem and Schur-product inequalities, we demonstrate theoretically that multiplicative coupling induces spectral contraction—yielding tighter eigenvalue bounds for the relevant matrix structures. This spectral contraction is known to improve optimization stability and efficiency. Consequently, schemes with this property are expected to learn the positional coupling function more rapidly and robustly in shallow layers, leading to its localization within a small set of attention heads. By contrast, additive mechanisms, lacking a direct content–relative position interaction in the logit, are predicted to underperform on tasks that critically depend on combining content with relative positional information.
To empirically validate our theoretical predictions, we design two synthetic tasks: Task 1, which requires content–relative positional reasoning, and Task 2, a position-agnostic control. We evaluate several PE methods (RoPE, Absolute, Relative (T5-style), ALiBi, NoPE) and conduct systematic layer- and head-wise ablations—zeroing out each head’s outputs—to probe the resulting internal structures. We also examine alternative coupling strategies, including MLA variants inspired by Deepseek-V3, to assess how different mechanisms and mixing timings shape positional processing.
Our results align closely with theory. On Task 1, requiring content-relative coupling, RoPE’s multiplicative mechanism consistently outperforms other methods, yielding superior generalization. Head-wise ablations confirm the predicted localization: RoPE produces an early-layer “single-head deposit,” wherein one shallow head carries the bulk of positional computations. By contrast, additive-bias schemes (Absolute, Relative, ALiBi) and NoPE exhibit diffuse or negligible head specialization in both tasks, and RoPE itself shows no deposit pattern on Task 2. Furthermore, MLA variants—specifically those employing a concatenation-based coupling mechanism (e.g., before the inner product)—produce a more distributed “average deposit” pattern. These findings provide strong empirical evidence that the form of content–position coupling fundamentally determines the model’s internal structure and learning dynamics.
  • We introduce a unified spectral-theoretic framework analyzing positional *interaction* in attention logits as structured (Toeplitz) matrices, distinguishing fundamental additive (bias/summation) and multiplicative (Hadamard) coupling mechanisms.
  • We show theoretically that multiplicative coupling (e.g., RoPE) induces spectral contraction—tighter eigenvalue bounds—which theory links to improved optimization stability. We empirically validate that PE methods employing these fundamental mechanisms exhibit predicted performance differences and deposit patterns (e.g., RoPE’s shallow single-head localization) on tasks requiring content-position coupling.
  • Via systematic layer- and head-wise ablations, we further investigate alternative strategies beyond fundamental mechanisms, such as MLA’s concatenation-based design. We demonstrate that these approaches yield contrasting, more distributed deposit patterns (e.g., average deposit vs. single-head localization), providing insights into how processing and combination choices influence learned internal structures.
The remainder of this paper is organized as follows. Section 3 details our spectral-theoretic framework. Section 4 describes the experimental setup and main results. Section 5 presents the ablation studies and mechanistic analysis.
2 Related Work
Positional Encoding in Transformers. Positional encoding was introduced alongside the Transformer architecture to endow self-attention with sequence order information vaswani2017attention . Early approaches employed learnable positional embeddings shaw2018self , while subsequent work integrated position biases directly into the attention logits, giving rise to relative position encoding and its extensions press2022alibi ; li2024fire ; chi-etal-2023-dissecting ; chi2022kerple ; raffel2020exploring ; SU2024127063 ; zhao2025drope . Among these, Rotary Positional Encoding (RoPE) SU2024127063 stands out by applying a multiplicative bias to the $QK^{\top}$ matrix, and has been widely adopted in modern open-source models. Related efforts on sliding-window and length-generalization designs are also often framed as positional encoding variants kiyono-etal-2021-shape ; chen2024clex ; chen2024longlora ; ding2024longrope ; peng2024yarn . Recent RoPE-based improvements include Frequency-Partitioned Encoding (FoPE) HuaFoPE2025 and the Multi-head Latent Attention (MLA) mechanism in Deepseek-V3 liu2024deepseek .
Mechanistic Analyses of Positional Encoding. Empirical studies have probed how Transformers internalize positional signals. Some works demonstrate that even without explicit positional embeddings (NoPE), models can learn positional distinctions wang-etal-2024-length ; zuo-etal-2025-position ; haviv-etal-2022-transformer ; barbero2024round , occasionally outperforming PE-equipped counterparts on select tasks kazemnejad2023impact . Explanations range from implicit causal masking zuo-etal-2025-position to dataset-dependent effects barbero2024round . Other analyses document characteristic behaviors of RoPE, such as “massive value” phenomena and rotation artifacts jinMassiveValuesSelfAttention2025 ; jonasson2025rotary .
Toeplitz Structure and Spectral Theory. The action of positional encoding in attention can be unified under Toeplitz-matrix representations bottcher2000analysis ; grenander2002toeplitz ; gray2006toeplitz ; oppenheim1999discrete . Classical results on Toeplitz spectral distributions, including Szegő’s theorem, inform our theoretical framework. Moreover, the impact of matrix spectra on optimization convergence is well studied in numerical linear algebra and convex optimization contexts horn1985matrix ; horn1991topics ; boyd2004convex ; nesterov2004introductory .
3 Theoretical Structure
3.1 Preliminaries and Assumptions
In Transformer architectures, the attention mechanism relies on the inner product between the query vector $q_{i}$ and the key vector $k_{j}$ at positions $i$ and $j$, respectively. This inner product, $q_{i}^{\top}k_{j}$, serves as the foundation for computing attention scores. Different positional encoding (PE) strategies influence this computation by injecting positional information in distinct ways.
Absolute Positional Encoding vaswani2017attention introduces position information at the input layer by directly adding a position vector $p_{i}$ to each token embedding $e_{i}$:
$h_{i}=e_{i}+p_{i}.$ (1)
Consequently, the attention scores are computed based on position-infused vectors:
$\alpha_{i,j}\propto\exp\bigl{(}(e_{i}+p_{i})^{\top}(e_{j}+p_{j})\bigr{)}.$ (2)
The residual connections throughout the Transformer layers ensure that this form of position-content coupling persists at deeper levels.
Relative Positional Encoding raffel2020exploring modifies the attention logits post inner-product, by adding a relative position-dependent bias term $b_{j-i}$:
$\alpha_{i,j}\propto\exp\bigl{(}q_{i}^{\top}k_{j}+b_{j-i}\bigr{)}.$ (3)
Here, $b_{j-i}$ is often a learnable parameter, either directly or as a function of relative distance $j-i$ parameterized through shared weights. This method allows position-dependent modifications without altering the token representations themselves.
Rotary Positional Encoding (RoPE) SU2024127063 instead applies a position-specific rotation matrix $R_{i}$ (or $R_{j}$) to each query and key vector before the inner product is computed. Due to the properties of these rotational transformations, the resulting dot product admits a key equivalence:
$q_{i}\leftarrow R_{i}q_{i},\quad k_{j}\leftarrow R_{j}k_{j}\quad\Longrightarrow\quad(R_{i}q_{i})^{\top}(R_{j}k_{j})=q_{i}^{\top}R_{j-i}k_{j}.$ (4)
Here, $R_{j-i}$ is a rotation matrix determined solely by the relative position difference. RoPE thus achieves a tightly coupled interaction between content and relative position within the attention score computation.
Notably, even in the absence of explicit positional encoding—commonly referred to as NoPE haviv-etal-2022-transformer —large Transformer models have demonstrated an ability to implicitly encode position through learned patterns in attention heads or depth-wise representation dynamics. This motivates our first key assumption:
Assumption 3.1 (Representation Decomposition).
Each token representation $x_{i}$ at position $i$ can be decomposed into a position-independent content component $c_{i}$ and a position-dependent component $p_{i}$, such that
$x_{i}=c_{i}+p_{i}.$ (5)
This decomposition provides a conceptual foundation to analyze how different PE mechanisms construct or interact with the positional component $p_{i}$, and how this interaction shapes attention.
To formalize the structural implications of positional interaction in attention, we invoke a classic object in matrix analysis:
Definition 3.1 (Toeplitz Matrix).
A matrix $A$ is a Toeplitz matrix if its entries are constant along each diagonal for some sequence ${a_{k}}$. An $N\times N$ Toeplitz matrix $T$ can be expressed as:
$T=\begin{bmatrix}a_{0}&a_{-1}&a_{-2}&\cdots&a_{-(N-1)}\\ a_{1}&a_{0}&a_{-1}&\cdots&a_{-(N-2)}\\ a_{2}&a_{1}&a_{0}&\cdots&a_{-(N-3)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{N-1}&a_{N-2}&a_{N-3}&\cdots&a_{0}\end{bmatrix},$ (6)
We denote this compactly as $T=[a_{i-j}]$.
Toeplitz matrices naturally emerge in systems with translation-invariant or shift-dependent structure, making them particularly suited to model attention patterns influenced by relative position.
Our second foundational assumption connects this matrix structure to positional information within the attention mechanism:
Assumption 3.2 (Toeplitz Structure from Positional Interaction).
The attention score matrix induced solely by the position-dependent components $p_{i}$ exhibits a Toeplitz structure:
$\mathbf{Attention}(p_{i},p_{j})=[a_{i-j}].$ (7)
Remark 3.1.
The equal sign is used here because the normalization operation does not change the Toeplitz property of the matrix.
Remark 3.2.
It can be verified that all forms of sine-cosine position embedding of absolute position encoding satisfy this assumption and the actual effect of relative position encoding is also presented in the form of a Toeplitz matrix.
Together, these two assumptions enable us to explore how positional information modulates the spectral behavior of attention matrices, and to quantify the effects of different PE strategies using tools from the spectral theory of Toeplitz operators.
3.2 Toeplitz Representation of Positional Encoding
In the Transformer’s attention mechanism, the interaction between representations at positions $i$ and $j$ is captured by the attention logit matrix $\mathbf{L}$, where $L_{i,j}=\text{AttentionLogit}_{i,j}$. Different positional encoding (PE) mechanisms influence this logit matrix in distinct ways to incorporate positional information.
Following our decomposition $x_{i}=c_{i}+p_{i}$, where $c_{i}$ is the content component and $p_{i}$ is the position component, and defining $q_{i}^{c}=W_{Q}c_{i}$, $q_{i}^{p}=W_{Q}p_{i}$, $k_{j}^{c}=W_{K}c_{j}$, and $k_{j}^{p}=W_{K}p_{j}$, we analyze the structure of $\mathbf{L}$.
We denote the Gram matrix of a set of column vectors $\{v_{i}\}$ and $\{w_{j}\}$ as $G_{v,w}$, where $(G_{v,w})_{i,j}=v_{i}^{\top}w_{j}$. Under this notation, for PE methods that operate via additive modification of token representations or attention biases (excluding RoPE), the attention logit matrix can be expressed as:
$\mathbf{L}_{\text{Non-RoPE}}=G_{q^{c},k^{c}}+G_{q^{c},k^{p}}+G_{q^{p},k^{c}}+G_{q^{p},k^{p}}+\mathbf{B}$ (8)
Here, $\mathbf{B}$ denotes an explicit bias matrix (e.g., from T5 or ALiBi). Under our theoretical assumptions, the term $G_{q^{p},k^{p}}$ forms a Toeplitz matrix, since it depends only on relative positions. Similarly, $\mathbf{B}$ is Toeplitz, as it typically encodes relative-position-dependent biases ($B_{i,j}=b_{i-j}$). The remaining terms involving content ($G_{q^{c},k^{c}}$, $G_{q^{c},k^{p}}$, $G_{q^{p},k^{c}}$) are not Toeplitz due to their dependence on specific token content.
In contrast, RoPE applies complex-valued rotations to token representations. By pairing adjacent dimensions into complex numbers, we construct complex-valued vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{p}_{i}$, and define a modulation vector $\mathbf{e}_{i}$ such that $\mathbf{e}_{i}^{(j)}=e^{i\theta_{j}}$. The logit matrix then takes the form:
$\mathbf{L}_{\text{RoPE\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ G_{\mathbf{e}}$ (9)
with $\circ$ denoting the Hadamard (element-wise) product, and the final logit is:
$\mathbf{L}_{\text{RoPE}}=\operatorname{Re}\left\{\mathbf{L}_{\text{RoPE\_complex}}\right\}$ (10)
Since $G_{\mathbf{e}}$ is Toeplitz by construction, the Hadamard product introduces relative-position dependence multiplicatively across all terms.
This formal structure allows us to compare how various PE mechanisms couple content and positional information. In the Non-RoPE case:
  • $G_{q^{c},k^{c}}$ encodes content-content similarity.
  • $G_{q^{p},k^{p}}$ captures position-position interactions and is Toeplitz.
  • $G_{q^{c},k^{p}}$ and $G_{q^{p},k^{c}}$ are cross terms encoding interactions between content and absolute position.
  • $\mathbf{B}$ encodes a content-independent, relative-position bias and is Toeplitz.
While $G_{q^{p},k^{p}}$ and $\mathbf{B}$ directly provide relative-position structure, they do not incorporate content-specific modulation. The cross-terms $G_{q^{c},k^{p}}$ and $G_{q^{p},k^{c}}$ may encode interactions between content and absolute position, but for these to collectively capture content-relative position coupling, the model would need to learn mappings where the cross-term interactions depend solely on $i-j$, yet are modulated by content. Such constrained mappings are difficult to realize in practice.
By contrast, RoPE’s multiplicative structure enables a more direct content-relative interaction: each component in the Gram matrix sum is modulated entrywise by the Toeplitz matrix $G_{\mathbf{e}}$, allowing relative-position structure to directly shape the contribution of content vectors. This results in a more expressive and learnable coupling mechanism between content and relative position.
Based on this analysis, we hypothesize that PE mechanisms enabling direct content-relative interaction within the attention logits will perform better on tasks requiring such coupling. RoPE, via its Hadamard product with $G_{\mathbf{e}}$, provides this mechanism, whereas Non-RoPE methods rely on structurally constrained additive approximations.
To test this hypothesis, we will design a synthetic evaluation task specifically tailored to probe the model’s ability to utilize content-relative positional information. This task will serve as our first benchmark for evaluating the expressive power of different positional encoding strategies.
3.3 Spectral Properties and Theoretical Prediction of Deposit Patterns
The structure of the attention logit matrix $\mathbf{L}$, particularly the properties of its Toeplitz or Toeplitz-like components, provides crucial insights into the stability and dynamics of positional information processing. The spectral properties of these matrices, especially the distribution and range of their eigenvalues, are known to influence the convergence of optimization algorithms.
For Toeplitz matrices, classical results such as Szegő’s theorem connect matrix structure to asymptotic eigenvalue distributions. We recall a relevant form below:
Theorem 3.1 (Szegő’s Theorem on Eigenvalue Distribution (From grenander2002toeplitz )).
Let $T_{N}(a)$ be an $N\times N$ Toeplitz matrix with $(T_{N}(a))_{i,j}=a_{i-j}$, and let $a(e^{i\theta})=\sum_{k=-\infty}^{\infty}a_{k}e^{ik\theta}$ be its symbol. If $a(e^{i\theta})$ is real-valued and continuous, then the eigenvalues $\lambda_{j}^{(N)}$ of $T_{N}(a)$ asymptotically fill the interval $[\min a(e^{i\theta}),\max a(e^{i\theta})]$ as $N\to\infty$.
Applying this theorem and related spectral analysis techniques to the Toeplitz structures induced by various PE mechanisms:
  • In Non-RoPE methods, the Toeplitz components $G_{q^{p},k^{p}}$ and $\mathbf{B}$ contribute additively to the logit, and their spectral ranges are determined by their respective coefficient sequences.
  • In RoPE, the complex logit involves a Hadamard product with the Toeplitz matrix $G_{\mathbf{e}}$. This multiplicative interaction theoretically contracts the eigenvalue range relative to additive compositions, resulting in tighter spectral bounds.
Spectral contraction—manifested as a reduced eigenvalue range or lower condition number—is well known to accelerate and stabilize gradient-based learning boyd2004convex ; gray2006toeplitz . The spectral properties associated with RoPE’s coupling mechanism, through this contraction, facilitate faster and more stable learning of the content-relative positional interaction. We give a more detailed analysis in Appendix B.
This enhanced convergence property leads to a concrete structural prediction: in early layers, the model tends towards specializing attention heads for content-relative positional coupling, while other heads in the same layer may exhibit less specific positional roles. This specialization results in a localized deposit pattern, where certain heads disproportionately contribute to processing positional information. Such behavior is expected to emerge more prominently under PE mechanisms with favorable spectral properties—most notably, RoPE—and to be absent or weaker in less structured encodings.
To validate this prediction, we design ablation experiments to confirm the emergence and characteristics of this deposit pattern.
4 Task Design and Experiment
4.1 Experiment Setup and Task Design
To evaluate our theoretical predictions on positional encoding and learning dynamics, we use a 6-layer Transformer decoder and compare six PE configurations: Absolute Positional Encoding, T5-style Relative Positional Encoding, ALiBi, NoPE (no positional encoding), Rotary Positional Encoding (RoPE), and random, fixed embeddings. Full training details—including hyperparameters, optimizer settings, and vocabulary construction—are provided in Appendix C.
We introduce two synthetic classification tasks that isolate positional requirements:
Task 1: Trigger Word Relative Distance Prediction. Each sequence contains two designated trigger words drawn from a small vocabulary (500–1000 tokens). The target label is the absolute index distance between these words. This task demands tight coupling of token identity and relative position. We generate 100,000 examples with a 7:3 train/test split.
Task 2: Trigger Word Counting. Each sequence is labeled by the total count of occurrences of a fixed set of 10–30 trigger words, irrespective of their positions. This position-agnostic task tests pure content recognition. We generate 50,000 examples with a 7:3 train/test split.
These tasks enable a controlled comparison of PE methods under position-sensitive (Task 1) and position-agnostic (Task 2) settings, directly testing our theoretical framework.
4.2 Convergence Analysis
We plot the experimental results as line graphs. The training accuracy curve indicates the convergence speed of each PE method on the training set, while the test accuracy curve reflects its capacity to learn the task’s underlying rules. Convergence results for Task 1 and Task 2 are displayed in Figures 2 and 3, respectively.
Figure 2: Training and test accuracy vs. epoch for different PE methods on Task 1 (relative-dis).
Task 1: Trigger Word Relative Distance Prediction. Leveraging its Hadamard-product coupling, RoPE converges fastest and most stably, reaching high training and test accuracy with a minimal generalization gap. By contrast, Absolute encoding ranks second: its fixed first-layer bias $G_{p}$ can be adjusted dynamically in deeper layers. Random initialization matches Absolute on training accuracy but overfits—its non-Toeplitz bias limits generalization, yielding lower test performance. Meanwhile, ALiBi performs poorly: its fixed, data-agnostic Toeplitz bias $B$ impedes adaptation and slows convergence. Neither NoPE nor standard Relative generalizes, since both lack a mechanism to bind content and relative position.
Figure 3: Training and test accuracy vs. epoch for different PE methods on Task 2 (counting).
Task 2: Trigger Word Count. In the position-agnostic Task 2, imposing positional biases generally degrades generalization. Absolute encoding incurs a mild penalty—its fixed $G_{p}$ can still be refined by later layers—while RoPE’s soft, data-adaptive bias imposes even less overhead, yielding superior test accuracy. Overfitting emerges under Random initialization: its non-Toeplitz bias perfectly fits content counts but produces the largest train–test gap. Relative encoding attenuates its learned bias toward zero and thus matches NoPE’s performance. ALiBi suffers most, as its rigid, fixed Toeplitz bias obstructs even simple counting.
5 Ablation Studies and Mechanistic Insights
5.1 Attention Head Specialization for Positional Processing
To complete our theoretical argument, we predict that PE mechanisms with spectrally stable position coupling—such as RoPE—will localize essential positional computations into a few specialized attention heads, especially in early layers, yielding a deposit pattern. We test this via a causal, head-wise ablation study: for each layer $l$ and head $h$, we zero out the head’s output $H_{l,h}$ and measure the resulting drop in Task 1 test accuracy relative to the intact model. This accuracy reduction quantifies each head’s criticality for content–position coupling.
Figure 4: Head-wise ablation violin plot for three situations.
Figure 4 confirms that under RoPE, ablating a single shallow-layer head causes a pronounced accuracy loss, indicating early-layer specialization. By contrast, no pronounced deposit pattern appears under No-RoPE methods like NoPE or ALiBi. In the main text we illustrate NoPE’s Task 1 ablation and RoPE’s Task 2 ablation in Figure 4; the full set of non-RoPE results (Absolute, ALiBi, Relative, Random) is provided in Appendix C.
Building on this finding, we hypothesize that only a subset of heads needs RoPE to capture relative position, while the remaining heads can operate without positional bias. We therefore conduct a “partial RoPE” experiment, applying RoPE to $k$ heads and NoPE to the rest. As Figure 5 shows, even with only a handful of RoPE-enabled heads, Task 1 performance remains below-near-optimal, and Task 2 generalization is shocked.
Figure 5: Performance when RoPE is applied to a subset of heads.
5.2 Modulating Deposit Patterns via MLA
Motivated by the observed deposit pattern and the “partial RoPE” results, we explore two strategies to diffuse positional responsibility across heads. First, we augment RoPE by adding an absolute positional embedding before the first layer, which provides an initial position cue that can be adjusted by deeper layers. Second, we adopt the MLA (Multi-head Latent Attention) mechanism from Deepseek-V3, which mixes NoPE and RoPE signals before the attention inner product.
We present the result of the early method in Figure 6. Due to the absolute position encoding, the deposit pattern no longer appears in the first layer. Ablating each head does not affect the overall performance. However, starting from the second layer, the deposit pattern reappears, which also indirectly reflects that the deposit pattern is an inherent phenomenon of the coupling between rope capture content and relative position. It will not be completely eliminated by adding absolute position information to the input.
Figure 6: (Left) The performance of RoPE with an absolute position on Task 1 and Task 2. (Right) "Deposit Pattern" of RoPE-Abs.
The late method MLA could solve the deposit pattern. Firstly, we use the inner integral calculation of an attention head to demonstrate MLA:
$<q_{i},k_{j}>=<(W_{UQ}W_{DQ}x_{i};R_{i}W_{UR}W_{DQ}x_{i}),(W_{UK}W_{DKV}x_{j};R_{j}W_{KR}x_{j})>,$ (11)
where (;)(\;;\;)( ; ) means the operation of vector stitching. Without considering the dimensionality reduction operation of MLA to improve computational efficiency, we can understand that each head independently processes the inner product of NoPE and RoPE once and sums them up, and the parameters of the two parts are independent and do not interfere with each other. In our theoretical frame work, it can be expressed as
$\mathbf{L}_{\text{MLA\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ(\alpha G_{\mathbf{e}}+I).$ (12)
We transplant this MLA module into our 6-layer decoder (see Appendix C for hyperparameters). Table 1 shows the MLA performance on the two tasks. Crucially, head-wise ablation under MLA (Figure 7) shows no single head is indispensable for Task 1, demonstrating effective mitigation of the earlier deposit pattern.
Figure 7: Head-wise ablation violin plot under MLA.
Method Task 1 Task 2
MLA 88.34 97.41
RoPE 92.64 69.43
NoPE 3.51 77.69
Table 1: MLA performance comparison on the two tasks.
Although RoPE was designed to enhance length generalization, real-world results often fall short of this ideal. We hypothesize that early single-head deposition hinders length robustness: when one head monopolizes positional processing, its representations may not effectively transfer to deeper layers under length shifts. The MLA mechanism, by diffusing positional responsibility, points toward a remedy. More broadly, our framework suggests that transforming the relative-position signal $G_{e}$ into a distributed form $F(G_{e})$—for example, via controlled mixing—could reconcile spectral stability with robust, length-agnostic encoding. We provide a specific qualitative statement of this view in Appendix D.
6 Conclusion and Limitations
In conclusion, we have introduced a spectral-theoretic framework that connects positional encoding designs to the eigenstructure of attention logits and shown—through synthetic tasks and head-wise ablations—that these spectral signatures reliably predict learning dynamics and internal processing patterns (e.g., RoPE’s “single-head deposit”). Our results illuminate how different coupling mechanisms shape the interplay between content and position in attention and suggest practical strategies—such as controlled mixing—to achieve both spectral stability and distributed encoding. These insights lay the groundwork for principled design of more robust and interpretable PE schemes in Transformer models.
Limitations: This study is primarily theoretical and relies on controlled synthetic tasks, so its findings may not directly translate to performance on complex, real-world datasets or broad benchmarks. We also explore only a limited range of task complexities and positional relationships; more intricate challenges—such as sequence reversal or Dyck-language recognition—likely exceed a Transformer’s positional reasoning capabilities. Finally, our proposed connection between single-head deposit patterns and length extrapolation remains theoretical and requires direct empirical validation.
Appendix A Notations List
The notations used throughout this article are summarized in Table 2.
Table 2: Some important notations used in this paper.
Notation Description
$x_{i}$ Real vector in $\mathbb{R}^{2n}$
$\mathbf{x}_{i}$ Complex vector in $\mathbb{R}^{n}$
$c$ Content label for a token or a hidden activation
$p$ Position label for a token or a hidden activation
$e_{i}$ Absolute position encoding for $i$
$b$ or $\mathbf{B}$ The bias decided by the relative position encoding
$R_{i}$ The rotary matrix for position $i$ of RoPE
$G_{u,v}$ The gram matrix for the vector columns $\{v_{i}\},\{u_{i}\}$
$\mathbf{L}$ Logit value of attention layer
$T_{N}(a)$ Toeplitz Matrix for $a_{i},i=-N,...,N$
Appendix B Detailed Theoretical Analysis
In this section of the appendix, we provide detailed theoretical analysis and derivations supporting the non-trivial conclusions presented in the main text. Our analysis relies on idealized proofs that simplify the complex optimization dynamics of neural networks, aiming to reveal a theoretical explanation for the observed experimental phenomena. While the actual simulation and analysis of non-convex optimization in full Transformer models is extremely difficult, this theoretical framework offers valuable insights into the underlying principles.
Our analysis is predicated on a core theoretical link regarding the role of spectral properties in learning. We rely on the established principle that favorable spectral properties of matrices involved in optimization—such as a more contracted eigenvalue spectrum (tighter bounds) or a smaller spectral norm—correlate with faster and more stable convergence of learning algorithms. Building upon this, our theoretical framework posits that the rapid and stable convergence of the positional coupling function, driven by these favorable spectral properties, will lead to a specific phenomenon during training: the localization of this function within a limited set of attention heads in shallow layers, resulting in the empirically observed deposit pattern [30].
Proposition B.1 (RoPE Logit Form): This proposition formally derives the mathematical form of the Rotary Positional Encoding (RoPE) attention logit as presented in the main text (Equation 9), showing how RoPE’s rotation mechanism translates into a specific inner product structure involving content and relative positional information. We would prove that
$\mathbf{L}_{\text{RoPE\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ G_{\mathbf{e}}.$
We use the fomula in [3] that a logit value of i and j in $\mathbf{L}_{\text{RoPE\_complex}}$ is
$\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)}e^{\mathrm{i}(i-j)\theta_{t}},$
and there are $t_{1}$ and $t_{2}$ that
$(\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)})e^{\mathrm{i}(i-j)\theta_{t_{1}}}\leq\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)}e^{\mathrm{i}(i-j)\theta_{t}}\leq(\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)})e^{\mathrm{i}(i-j)\theta_{t_{2}}},$
Due to continuity and the mean value theorem, there exists $t_{i-j}$ that
$\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)}e^{\mathrm{i}(i-j)\theta_{t}}=(\sum_{k=0}^{d/2-1}\mathbf{q}_{i}^{(t)}\mathbf{k_{j}}^{(t)})e^{\mathrm{i}(i-j)\theta_{t_{i-j}}}=\mathbf{q_{i}}\mathbf{k_{j}}e^{\mathrm{i}(i-j)\theta_{t_{i-j}}}.$
Then we can use $e^{\mathrm{i}(i-j)\theta_{t_{i-j}}}$ to generate $G_{e}$ to finish our proof.
Proposition B.2 (Logit Expressiveness): This proposition analyzes and demonstrates the higher degree of freedom and expressive power of the RoPE attention logit form (Equation 9) compared to the non-RoPE additive logit form (Equation 8). It highlights why RoPE’s structure is theoretically better suited to capture the complex content-relative positional relationships required by tasks like Task 1.
Task 1 is directly related to the following optimization problem:
For $-i\leq s\leq d-s$ and constant $C_{1}$ and $C_{2}$,
1. Relative Position Encoding
$<q_{i},k_{j}>+b_{i-j}>C_{1}>C_{2}><q_{i+s},k_{j+s}>+b_{i-j},$
2. RoPE
$\mathbf{Re}\{<\mathbf{q}_{i},\mathbf{k}_{j}>e^{\mathrm{i}(i-j)\theta{t_{i-j}}}\}>C_{1}>C_{2}>\mathbf{Re}\{<\mathbf{q}_{i+s},\mathbf{k}_{j+s}>e^{\mathrm{i}(i-j)\theta{t_{i-j}}}\}.$
It is easy to verify that the solution space of RoPE is much larger than that of Relative Position Encoding. This can be considered from the following two aspects. First, the expression of RoPE can make the key-value pairs with a fixed distance meet the requirements as long as the angle distribution meets the requirements, while the expression of Relative Position Encoding requires a more stringent angle distribution because the bias has no significant effect. Second, RoPE actually only needs a significant angle to easily meet the conditions, which actually corresponds to the cause of Massive Value described in  [22].
Proposition B.3 (Spectral Contraction): This proposition formally proves the spectral property claimed in the main text. By applying theorems like Szegő’s theorem and inequalities concerning the Hadamard product, it demonstrates that the Toeplitz matrix structure associated with multiplicative coupling (as in RoPE) has a more desirable eigenvalue spectrum (e.g., more compact or tighter bounds) compared to Toeplitz structures associated with additive coupling mechanisms.
We simplify the proposition into the following: Let $W,E$ be $N\times N$ positive definite Toeplitz matrices. Then the eigenvalue spectrum of the Hadamard product $W\circ E$ is more compact than $W$.
Here, each element of E is a complex number with a membrane length of 1. We can directly use Schur’s inequality to prove that the membrane length of the eigenvalue of $W\circ E$ must be less than or equal to the membrane length of the eigenvalue of $W$.
However, a strict comparison requires citation of Szegő’s Theorem on Eigenvalue Distribution [25]. So let us calculate the generating function (sign function) corresponding to $W\circ E$ and $W$ respectively.
$\mathbf{Symble}(W\circ E)(\theta)=\sum_{i=-N+1}^{N-1}w_{i}e^{\mathrm{i}(i\theta_{i}+i\theta)}=\sum_{i=0}^{N-1}2\mathbf{Re}\{w_{i}\}\cos(i\theta_{i}+i\theta)$
$\mathbf{Symble}(W)(\theta)=\sum_{i=-N+1}^{N-1}w_{i}e^{\mathrm{i}(i\theta)}=\sum_{i=0}^{N-1}2\mathbf{Re}\{w_{i}\}cos(i\theta)$
We will use the following lemma to prove that the range of the above formula must be smaller than that of the following formula, and thus Szegő’s Theorem shows that its eigenvalue spectrum is more compact.
Lemma B.1 (Amplitude bound and phase–alignment criterion).
Fix an integer $N\geq 1$ and non-negative weights $w_{i}\geq 0$. Define
$f(\theta)\;:=\;\sum_{i=0}^{N-1}2w_{i}\cos\!\bigl{(}i\theta_{i}+i\theta\bigr{)},\qquad g(\theta)\;:=\;\sum_{i=0}^{N-1}2w_{i}\cos(i\theta),\qquad\theta\in\mathbb{R}.$
Then
[…]
Equality holds iff all phase offsets coincide modulo $2\pi$:
$\theta_{0}\equiv\theta_{1}\equiv\cdots\equiv\theta_{N-1}\pmod{2\pi}.$
Consequently the peak-to-peak range of $g$ (width $4\sum_{i}w_{i}$) never falls below that of $f$, and both ranges coincide only under complete phase alignment.
Proof.
Denote
$P(z):=\sum_{i=0}^{N-1}c_{i}z^{\,i},\quad c_{i}:=w_{i}e^{\mathrm{i}\theta_{i}},\qquad Q(z):=\sum_{i=0}^{N-1}w_{i}z^{\,i},\quad|z|=1.$
Upper bound.
For any $|z|=1$ the triangle inequality gives
$|P(z)|\;\leq\;\sum_{i=0}^{N-1}|c_{i}|\;=\;\sum_{i=0}^{N-1}w_{i}.$
Hence
[…]
Conversely, $\bigl{|}\cos(i\theta)\bigr{|}\leq 1$ implies $|g(\theta)|\leq 2\sum_{i}w_{i}$ for every $\theta$, while $g(0)=2\sum_{i}w_{i}$ shows that this bound is attained, so $\max_{\theta}|g(\theta)|=2\sum_{i}w_{i}$. Combining the two yields the desired inequality.
Equality case.
Suppose $\max_{\theta}|f(\theta)|=2\sum_{i}w_{i}$. Then at some $\theta^{\star}$ both previous inequalities are tight:
$\bigl{|}P(e^{\mathrm{i}\theta^{\star}})\bigr{|}=\sum_{i}w_{i},\qquad\bigl{|}\Re P(e^{\mathrm{i}\theta^{\star}})\bigr{|}=\bigl{|}P(e^{\mathrm{i}\theta^{\star}})\bigr{|}.$
The first equality forces **all terms in the sum for $P$ to share a common phase**, i.e. $e^{\mathrm{i}(\theta_{i}+i\theta^{\star})}$ is the same for every $i$; the second forces this common phase to be 00 or $\pi$, which does not affect alignment. Thus
$\theta_{i}+i\theta^{\star}\equiv\theta_{0}\pmod{2\pi}\quad\forall i\;\Longrightarrow\;\theta_{i}\equiv\theta_{0}\pmod{2\pi}\quad\forall i.$
Conversely, if all $\theta_{i}$ are equal, choosing $\theta=-\theta_{0}$ gives $f(-\theta_{0})=g(0)=2\sum_{i}w_{i}$, so equality is achieved.
The lemma follows. ∎
Appendix C Support Experiment
C.1 Experiment Setup
All experiments were performed on a single NVIDIA RTX 4090.
The experimental parameter settings of non-MLA are shown in Table 3, and the experimental parameter settings of MLA are shown in Table 4.
Setting Value
Vocabulary size (vocab_size) 574
Model dimension ($d_{\mathrm{model}}$) 256
Feed-forward dimension ($d_{\mathrm{ff}}$) 512
Number of attention heads (num_heads) 16
Number of decoder layers (num_layers) 6
Dropout rate 0.1
Maximum sequence length (max_len) 128
Positional encoding types tested nope, absolute, alibi, relative, random, rope
Number of epochs 100 (Task 1);150 (Task 2)
Batch size 512 or 1024
Optimizer AdamW (lr=1e-4, betas=(0.98, 0.9), weight_decay=1e-5)
Learning-rate scheduler Cosine schedule with warm-up (6% of total steps)
Loss function Cross-entropy
Train/Test split 70% / 30%
Random seed 42
Table 3: Hyperparameters and settings (PE).
Setting Value
Vocabulary size (vocab_size) 574
Model hidden dimension ($d_{\mathrm{model}}$) 256
Feed-forward dimension ($d_{\mathrm{ff}}$) 512
Number of logical MLA heads (num_heads) 16 (8 in fact)
Compression dimension ($d_{\mathrm{compress}}$) 128
Number of decoder layers (num_layers) 6
Dropout rate 0.1
Maximum sequence length (max_len) 128
Positional embedding Rotary (RoPE)
Number of epochs 100 (Task 1); 150 (Task 2)
Batch size 1024
Optimizer AdamW (lr=1e-4, betas=(0.98, 0.9), weight_decay=1e-5)
Learning-rate scheduler Cosine warm-up (6% of total steps)
Loss function Cross-entropy
Train/Test split 70% / 30%
Random seed 42
Table 4: Hyperparameters and settings (MLA).
Remark C.1.
Note that all our experiments used dropout, so the deposit pattern is not a phenomenon that can be improved by ordinary dropout.
Since Task 2 did not have a position effect, our subsequent experiments were all on Task 1.
C.2 Experiments of Head Ablation on Other PE Methods
In this subsection, we use violin plots to show the ablation experiment results of Alibi, Absolute PE, and Relative PE.
In Figure 8, we can find that the position information of relative position encoding and ALiBi are scattered, while the absolute position encoding has some significant heads and some insignificant heads in the shallow layer, and the whole is oscillating.
Figure 8: Ablation experiments on other PE methods.
The similarities among NoPE, Relative PE, and ALiBi are consistent with previous studies on how NoPE obtains position information through causal masks [5, 20].
C.3 Some Experiments on Partial RoPE
It is foreseeable that MLA can fully alleviate the deposit pattern at the expense of a little generalization ability. We also conducted some ablation experiments in the experiments of RoPE for some heads and NoPE for some heads. Interestingly, we found that the most significant deposit pattern is not necessarily in the first layer, but may be in the second or third layer. In some cases, the first layer does not even have a deposit pattern. We can imagine that this is the effect of using NoPE attention heads in deeper layers. This inspires us that the way the FFN affects the position encoding is far more complicated than we imagined. We give the partial ablation line graph of the first k heads doing RoPE layer l below.
Figure 9: $k=2$, $l=1,2$.
Figure 10: $k=3$, $l=1,2$.
Figure 11: $k=5$, $l=1,3$.
Figure 12: $k=6$, $l=1,2$.
Figure 13: $k=13$, $l=1,2$.
Figure 14: $k=15$, $l=2,4$.
C.4 Why Position Embedding Complex?
We conclude this section by explaining why understanding positional encodings is difficult and almost impossible to decouple.
We first give our conclusion that the position component in the token comes from manifold embedding rather than subspace decomposition, which means that extracting the position component requires us to have an algorithm to solve the changes in the position manifold. In an extremely ideal case, we can use the Levinson–Durbin algorithm in signal processing to recursively solve it through the value of the Toeplitz matrix. However, extracting a Toeplitz matrix from a general matrix is not unique, and we cannot determine whether the extracted Toeplitz matrix truly represents the position information.
This conclusion comes from a very simple observation: the position embedding of absolute position encoding is not orthogonal to word embedding. Therefore, there is no subspace decomposition that can completely separate the two components.
If we denote f as a continuous embedding from $S_{1}\otimes S_{1}...\otimes S_{1}$ to $\mathbb{R}^{n}$, and g as a continuous embedding from $\mathbb{R}_{+}$ to $S_{1}\otimes S_{1}...\otimes S_{1}$, then the essence of absolute position encoding is actually $f(g(i))$. In other words, the explicit position encoding that we can understand is essentially a manifold embedding.
The general method of dealing with manifold embeddings requires a specific embedding mapping. For general positional encodings, this mapping is not solvable, so it is almost impossible to decouple the positional encoding.
Appendix D Relationship to Length Generation
Here, we give a more detailed qualitative statement to describe our thinking on the relationship between deposit patterns and length generation.
More than one study has shown that transformer’s understanding of position information starts from the first layer [19, 9, 5]. The key to successful length generalization is that responses to relative position information within the training length can be naturally transferred to a longer length (test length). Therefore, the lower the coupling between location and content, the better the scalability. If the coupling degree is high, it is necessary to learn the coupling mode.
Our experimental results show that the deposit pattern is not difficult to improve through dropout, which means that once a head of trasnformer of RoPE captures the pattern of training position information, the other heads no longer capture the position information. Therefore, the number of parameters responding to position content coupling is much less than that of relative position encoding and no position encoding, resulting in higher compression and lower regularization.
At the same time, Deposit Patterns shows that only some frequency bands are fully trained during training like [4]. These frequency bands are passed through the fully connected layer, spreading the influence to most of the frequency bands in the deep layer, which means that in fact only specific relative position differences are considered by the training process, so the length generalization ability is drastically affected.